OpenStudy (anonymous):

Model the function rule y = -2x - 3 with a table of values. I can show you my answer if you would like to tell me what I did wrong..

jimthompson5910 (jim_thompson5910):

Sure post what you have so far

OpenStudy (anonymous):

0 -1.5 2 - 2.5 4 -3.5 6 -4.5 8 -5.5

jimthompson5910 (jim_thompson5910):

when x = 0, y is .... y = -2x - 3 y = -2*0 - 3 ... replace x with 0 (since x = 0) y = 0 - 3 y = -3 So when x = 0, y = -3 so the first row should be: 0 -3

jimthompson5910 (jim_thompson5910):

similarly, when x = 2, y is... y = -2x - 3 y = -2*2 - 3 y = -4 - 3 y = -7 so the next row is: 2 -7 hopefully this is making sense

OpenStudy (anonymous):

Can you help me witht the next question? I suck at these..

jimthompson5910 (jim_thompson5910):

just keep practicing and you'll get better

jimthompson5910 (jim_thompson5910):

what's your next question

OpenStudy (anonymous):

For the function rule y = -2x - 3, provide the x-y coordinates of the y-intercept and the x-intercept

OpenStudy (anonymous):

so do i just put down what I got for this question?

jimthompson5910 (jim_thompson5910):

the y-intercept happens when x = 0 so we already know what the y-intercept is (look above)

jimthompson5910 (jim_thompson5910):

what is the y-intercept?

OpenStudy (anonymous):

- 3?

jimthompson5910 (jim_thompson5910):

as an ordered pair though

OpenStudy (anonymous):

(0,-3)

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

how would you find the x-intercept?

OpenStudy (anonymous):

Constant rate of change?

jimthompson5910 (jim_thompson5910):

no, you plug in y = 0 and solve for x

OpenStudy (anonymous):

x- y Coordinates (0,-3),(2,-7),(4,-11),(6,-15),(8,-19)

jimthompson5910 (jim_thompson5910):

yes those are points that lie on the line

jimthompson5910 (jim_thompson5910):

what about the x-intercept?

OpenStudy (anonymous):

I don't know.. :(

jimthompson5910 (jim_thompson5910):

plug in y = 0 and solve for x

jimthompson5910 (jim_thompson5910):

y = -2x - 3 0 = -2x - 3 x = ??

OpenStudy (anonymous):

-2x?

jimthompson5910 (jim_thompson5910):

0 = -2x - 3 3 = -2x -2x = 3 x = ???

OpenStudy (anonymous):

x = 0?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

try it again

OpenStudy (anonymous):

i don't know. I give up.

jimthompson5910 (jim_thompson5910):

-2 * _______ = 3 what goes in the blank?

jimthompson5910 (jim_thompson5910):

it's not 0

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

I really don't know. I'm doing homeschooling and i'm having trouble.

jimthompson5910 (jim_thompson5910):

what operation is the opposite of multiplication?

OpenStudy (anonymous):

Dividing

jimthompson5910 (jim_thompson5910):

so you divide both sides by _____ (fill in the blank) to isolate x -2x = 3 x = ???